OpenStudy (anonymous):

(sec(x)+tan(x))((1-sin(x))/cos(x))=1

OpenStudy (joannablackwelder):

Use sinx/cosx=tanx

OpenStudy (joannablackwelder):

And secx=1/cosx

OpenStudy (lynfran):

\[(\frac{ 1 }{ cosx }+\frac{ sinx }{ cosx })(\frac{ 1-sinx }{ cosx })=1\]\[(\frac{ 1+sinx }{ cosx })(\frac{ 1-sinx }{ cosx })=1\]\[\frac{ 1-\sin ^{2}x }{ \cos ^{2}x }=1\]...now use the Pythagorean identity to simplify

OpenStudy (anonymous):

i have already done that. now what

OpenStudy (joannablackwelder):

Solve the pythagorean identity for cos^2x and see what you get.

OpenStudy (anonymous):

how do you use the Pythagorean identity for cos^2x

OpenStudy (joannablackwelder):